# Long-time asymptotics of (1,3)-sign solitary waves for the damped nonlinear Klein-Gordon equation

### v1

@@ 5.4. Rough estimates for lengths and angles
In this subsection, using Lemma 5.7, we prove that $c_{12},\ c_{13},\ c_{23}$ converge to $\frac{1}{2}$. To this end, we first estimate the sum of the remaining two inner products in terms of the maximal one, keeping (3.9) in mind.
@@
$\begin{aligned} &\quad 1+\mathcal{N}-\left(\max{(c_{12},c_{13},c_{23})}-1\right)^{2}-\mbox{{mid}}{(c_{12},c_{13},c_{23})}\min{(c_{12},c_{13},c_{23})}\\ &\geq\frac{2}{3}\max{\left(c_{12}+\frac{1}{2},c_{13}+\frac{1}{2},c_{23}+\frac{1}{2}\right)}.\end{aligned}$ (5.48)
@@
The right inequality in (5.49) is obvious, so it suffices to prove the left one. This is clear when $c_{12}=1$. When $c_{12}<1$, it follows from
@@
Letting $x=c_{13}+c_{23}$, we have $\sqrt{2(c_{12}+1)}\leq x\leq 2c_{12}$, and in this case (5.48) becomes
$\quad 1+\mathcal{N}-\left(\max{(c_{12},c_{13},c_{23})}-1\right)^{2}-\mbox{{mid}}{(c_{12},c_{13},c_{23})}\min{(c_{12},c_{13},c_{23})}$
@@
$\quad-x^{2}+(2-c_{12})x+6-2c_{12}^{2}+4c_{12}$
$=-2(c_{12}+1)-(2-c_{12})\sqrt{2(c_{12}+1)}+6-2c_{12}^{2}+4c_{12}$
$=-\nu_{12}^{2}-\left(3-\frac{\nu_{12}^{2}}{2}\right)\nu_{12}+6-\frac{(\nu_{12}^{2}-2)^{2}}{2}+2\nu_{12}^{2}-4$
$=\frac{\nu_{12}(\nu_{12}-1)(6-\nu_{12}^{2})}{2}.$
In particular, by (3.9) and $c_{12}=\max{(c_{12},c_{13},c_{23})}$, we have $c_{12}\geq-\frac{1}{2}$, hence $\nu_{12}\geq 1$, and therefore
$\frac{\nu_{12}(\nu_{12}-1)(6-\nu_{12}^{2})}{2}$ $\geq\nu_{12}-1$
$=\frac{2c_{12}+1}{\sqrt{2(c_{12}+1)}+1}$
$\geq\frac{2}{3}\left(c_{12}+\frac{1}{2}\right).$
Thus, we have
@@
Hence (5.48) holds in the case $c_{23}\leq c_{13}\leq c_{12}$, and the lemma follows. ∎
Using this lemma, we estimate $c_{12},\ c_{13},\ c_{23}$.
@@
Let $\vec{u}$ be a (1,3)-sign solitary waves. Then, we have
$\left|c_{12}(t)+\frac{1}{2}\right|+\left|c_{23}(t)+\frac{1}{2}\right|+\left|c_{13}(t)+\frac{1}{2}\right|\lesssim\frac{1}{(\log{t})^{\frac{1}{2}}}$ (5.53)
@@
First, by Lemma 5.7, there exists $C>0$ such that for $t\geq 0$,
@@
Here, we assume that there exists $t^{\prime}>0$ such that
$\max{\left(c_{12}(t^{\prime})+\frac{1}{2},c_{13}(t^{\prime})+\frac{1}{2},c_{23}(t^{\prime})+\frac{1}{2}\right)}>\frac{45C}{L^{\frac{1}{2}}}.$ (5.57)
@@
$\max{\left(c_{12}(t)+\frac{1}{2},c_{13}(t)+\frac{1}{2},c_{23}(t)+\frac{1}{2}\right)}\geq\frac{45C}{L^{\frac{1}{2}}}.$ (5.58)
@@
Since $L^{-\frac{1}{2}}$ is decreasing, (5.58) continues to hold for $0<t-T_{1}\ll 1$, contradicting the maximality of $T_{1}$. Therefore $T_{1}=\infty$. In particular, the above computation shows that $\max(c_{12},c_{13},c_{23})$ is monotone increasing, and since it is bounded above by (5.56), it converges as $t\to\infty$ to some $c_{\max}\in\left(-\frac{1}{2},\frac{1}{3}\right]$. For $t$ sufficiently large, if $c_{12}$ is the maximum, then
@@
From this, taking $T_{3}>T_{2}\gg 1$ sufficiently large, we obtain
$1$ $\gtrsim\max{(c_{12}(T_{3}),c_{13}(T_{3}),c_{23}(T_{3}))}-\max{(c_{12}(T_{2}),c_{13}(T_{2}),c_{23}(T_{2}))}$
$\gtrsim\int_{T_{2}}^{T_{3}}\frac{ds}{s\log{s}}.$
For $T_{3}\gg T_{2}\gg 1$, this is impossible, which is a contradiction. Hence the lemma is proved.
@@ 6.1. Sharp estimates for lengths and angles
From the previous section, we obtained (5.53). This allows us to refine the estimates for $\rho_{1},\rho_{2},\rho_{3}$.
@@
Let $\vec{u}$ be a (1,3)-sign solitary waves. Then, we have
@@
We first estimate $a_{1},a_{2},a_{3}$ in terms of $d_{12},d_{13},d_{23}$. Before the proof, we set
@@
Lemma 6.2.
Let $\vec{u}$ be a (1,3)-sign solitary waves. Then, we have
$|a_{1}|+|a_{2}|+|a_{3}|+|d_{12}|+|d_{13}|+|d_{23}|\lesssim\frac{1}{(\log{t})^{4}}.$ (6.5)
@@
$\begin{aligned} \left|a_{1}-\frac{1}{5}(4d_{12}+4d_{13}+2d_{23})\right|&\lesssim\frac{|d|}{(\log{t})^{\frac{1}{4}}}+\frac{1}{(\log{t})^{4}},\\ \left|a_{2}-\frac{1}{5}(4d_{12}+2d_{13}+4d_{23})\right|&\lesssim\frac{|d|}{(\log{t})^{\frac{1}{4}}}+\frac{1}{(\log{t})^{4}},\\ \left|a_{3}-\frac{1}{5}(2d_{12}+4d_{13}+4d_{23})\right|&\lesssim\frac{|d|}{(\log{t})^{\frac{1}{4}}}+\frac{1}{(\log{t})^{4}}.\end{aligned}$ (6.6)
We first refine the estimates for the differential equations satisfied by $a_{1},a_{2},a_{3}$. By (2.12), we have
$\left|\mathcal{F}(\rho_{k})-(1-a_{k})\mathcal{F}(L)\right|\lesssim\frac{\mathcal{F}(L)|a|}{L^{\frac{1}{2}}}.$ (6.7)
@@
$\quad+O\left(\frac{|a|}{t(\log{t})^{\frac{1}{2}}}+\frac{1}{t(\log{t})^{8}}\right),$
@@
$\quad+O\left(\frac{|a|}{t(\log{t})^{\frac{1}{2}}}+\frac{1}{t(\log{t})^{8}}\right).$
By the same calculation, $a_{2},a_{3}$ satisfy
$\begin{aligned} \dot{a}_{1}&=\left(-2a_{1}+\frac{1}{2}a_{2}+\frac{1}{2}a_{3}+d_{12}+d_{13}\right)\mathcal{F}(L)\\ &\quad+O\left(\frac{|a|}{t(\log{t})^{\frac{1}{2}}}+\frac{1}{t(\log{t})^{8}}\right),\\ \dot{a}_{2}&=\left(\frac{1}{2}a_{1}-2a_{2}+\frac{1}{2}a_{3}+d_{12}+d_{23}\right)\mathcal{F}(L)\\ &\quad+O\left(\frac{|a|}{t(\log{t})^{\frac{1}{2}}}+\frac{1}{t(\log{t})^{8}}\right),\\ \dot{a}_{3}&=\left(\frac{1}{2}a_{1}+\frac{1}{2}a_{2}-2a_{3}+d_{13}+d_{23}\right)\mathcal{F}(L)\\ &\quad+O\left(\frac{|a|}{t(\log{t})^{\frac{1}{2}}}+\frac{1}{t(\log{t})^{8}}\right).\end{aligned}$ (6.8)
@@
$\quad+O\left(\frac{|a|}{t(\log{t})^{\frac{3}{2}}}+\frac{1}{t(\log{t})^{9}}\right).$
@@
$\begin{aligned} \dot{d}_{12}&=\left(3d_{12}+\frac{3}{2}d_{13}+\frac{3}{2}d_{23}-\frac{3}{4}a_{1}-\frac{3}{4}a_{2}+\frac{3}{2}a_{3}\right)\frac{\mathcal{F}(L)}{L}\\ &\quad+O\left(\frac{|a|+|d|}{t(\log{t})^{\frac{3}{2}}}+\frac{1}{t(\log{t})^{9}}\right),\\ \dot{d}_{13}&=\left(\frac{3}{2}d_{12}+3d_{13}+\frac{3}{2}d_{23}-\frac{3}{4}a_{1}+\frac{3}{2}a_{2}-\frac{3}{4}a_{3}\right)\frac{\mathcal{F}(L)}{L}\\ &\quad+O\left(\frac{|a|+|d|}{t(\log{t})^{\frac{3}{2}}}+\frac{1}{t(\log{t})^{9}}\right),\\ \dot{d}_{23}&=\left(\frac{3}{2}d_{12}+\frac{3}{2}d_{13}+3d_{23}+\frac{3}{2}a_{1}-\frac{3}{4}a_{2}-\frac{3}{4}a_{3}\right)\frac{\mathcal{F}(L)}{L}\\ &\quad+O\left(\frac{|a|+|d|}{t(\log{t})^{\frac{3}{2}}}+\frac{1}{t(\log{t})^{9}}\right).\end{aligned}$ (6.9)
@@
$|\dot{d}_{12}|+|\dot{d}_{13}|+|\dot{d}_{23}|\lesssim\frac{|a|+|d|}{t\log{t}}+\frac{1}{t(\log{t})^{9}}.$ (6.10)
@@
Then, by direct computation. we have
@@
$\begin{aligned} \dot{\xi}_{1}&=\left(-2\xi_{1}+\frac{1}{2}\xi_{2}+\frac{1}{2}\xi_{3}\right)\mathcal{F}(L)+O\left(\frac{|\xi|+|d|}{t(\log{t})^{\frac{1}{2}}}+\frac{1}{t(\log{t})^{8}}\right),\\ \dot{\xi}_{2}&=\left(\frac{1}{2}\xi_{1}-2\xi_{2}+\frac{1}{2}\xi_{3}\right)\mathcal{F}(L)+O\left(\frac{|\xi|+|d|}{t(\log{t})^{\frac{1}{2}}}+\frac{1}{t(\log{t})^{8}}\right),\\ \dot{\xi}_{3}&=\left(\frac{1}{2}\xi_{1}+\frac{1}{2}\xi_{2}-2\xi_{3}\right)\mathcal{F}(L)+O\left(\frac{|\xi|+|d|}{t(\log{t})^{\frac{1}{2}}}+\frac{1}{t(\log{t})^{8}}\right).\end{aligned}$ (6.13)
@@
$\quad+O\left(\frac{|\xi|^{2}+|d|^{2}}{t(\log{t})^{\frac{1}{2}}}+\frac{1}{t(\log{t})^{8}}\right).$
Thus, there exists $c_{1}>0$ such that
$\frac{d}{dt}|\xi|^{2}\leq-\frac{c_{1}}{t}|\xi|^{2}+\frac{1}{c_{1}}\left(\frac{|d|^{2}}{t(\log{t})^{\frac{1}{2}}}+\frac{1}{t(\log{t})^{8}}\right).$ (6.14)
By Gronwall’s inequality, we obtain
$|\xi|^{2}\lesssim\frac{|d|^{2}}{(\log{t})^{\frac{1}{2}}}+\frac{1}{(\log{t})^{8}},$
@@
By (6.6). and (6.9), there exists $c_{2}>0$ such that
$\begin{aligned} \left|\dot{d}_{12}-\left(\frac{12}{5}d_{12}+\frac{9}{5}d_{13}+\frac{9}{5}d_{23}\right)\frac{\mathcal{F}(L)}{L}\right|&\leq c_{2}\left(\frac{\mathcal{F}(L)}{L^{\frac{5}{4}}}|d|+\frac{\mathcal{F}(L)}{L^{5}}\right),\\ \left|\dot{d}_{13}-\left(\frac{9}{5}d_{12}+\frac{12}{5}d_{13}+\frac{9}{5}d_{23}\right)\frac{\mathcal{F}(L)}{L}\right|&\leq c_{2}\left(\frac{\mathcal{F}(L)}{L^{\frac{5}{4}}}|d|+\frac{\mathcal{F}(L)}{L^{5}}\right),\\ \left|\dot{d}_{23}-\left(\frac{9}{5}d_{12}+\frac{9}{5}d_{13}+\frac{12}{5}d_{23}\right)\frac{\mathcal{F}(L)}{L}\right|&\leq c_{2}\left(\frac{\mathcal{F}(L)}{L^{\frac{5}{4}}}|d|+\frac{\mathcal{F}(L)}{L^{5}}\right).\end{aligned}$ (6.15)
@@
Then, by considering the possible sign patterns among $d_{12},d_{13},d_{23}$ (i.e., the numbers of positive and negative entries), we obtain
$\frac{1}{4}|d|\leq\max{(d_{12},d_{13},d_{23})}\leq 3|d|.$ (6.17)
@@
$\max{(d_{12}(t^{\prime}),d_{13}(t^{\prime}),d_{23}(t^{\prime}))}>\frac{100c_{2}}{L^{4}(t^{\prime})}.$ (6.18)
@@
Furthermore, we assume that $T_{1}<\infty$. Then we assume
@@
Then, by (6.15) we have at $t=T_{1}$
$\dot{d}_{12}$ $\geq\left(\frac{12}{5}d_{12}+\frac{9}{5}d_{13}+\frac{9}{5}d_{23}\right)\frac{\mathcal{F}(L)}{L}-c_{2}\left(\frac{\mathcal{F}(L)}{L^{\frac{5}{4}}}|d|+\frac{\mathcal{F}(L)}{L^{5}}\right)$
$\geq\frac{3}{5}\frac{\mathcal{F}(L)}{L}d_{12}-c_{2}\left(\frac{\mathcal{F}(L)}{L^{\frac{5}{4}}}|d|+\frac{\mathcal{F}(L)}{L^{5}}\right)$

### v2

@@ 5.4. Rough estimates for lengths and angles
In this subsection, using Lemma 5.6, we prove that $c_{12},\ c_{13},\ c_{23}$ converge to $\frac{1}{2}$ as $t\to\infty$. To this end, we first control the sum of the other two inner products by the largest one.
@@
$\begin{aligned} &\quad\mathcal{N}-\left(\max{(c_{12},c_{13},c_{23})}\right)^{2}+2\max{(c_{12},c_{13},c_{23})}\\ &\quad-\mbox{{mid}}{(c_{12},c_{13},c_{23})}\min{(c_{12},c_{13},c_{23})}\\ &\geq\frac{2}{3}\max{\left(c_{12}+\frac{1}{2},c_{13}+\frac{1}{2},c_{23}+\frac{1}{2}\right)}.\end{aligned}$ (5.48)
@@
The right inequality in (5.49) is obvious, so it suffices to prove the left one. When $c_{12}=1$, it is obvious. When $c_{12}<1$, it follows from
@@
In this proof, we define $x=c_{13}+c_{23}$. Then, we have $\sqrt{2(c_{12}+1)}\leq x\leq 2c_{12}$, and (5.48) becomes
$\quad\mathcal{N}-\left(\max{(c_{12},c_{13},c_{23})}\right)^{2}+2\max{(c_{12},c_{13},c_{23})}$
$\quad-\mbox{{mid}}{(c_{12},c_{13},c_{23})}\min{(c_{12},c_{13},c_{23})}$
@@
$\begin{aligned} &\quad-x^{2}+(2-c_{12})x+6-2c_{12}^{2}+4c_{12}\\ &=-2(c_{12}+1)-(2-c_{12})\sqrt{2(c_{12}+1)}+6-2c_{12}^{2}+4c_{12}\\ &=-\nu_{12}^{2}-\left(3-\frac{\nu_{12}^{2}}{2}\right)\nu_{12}+6-\frac{(\nu_{12}^{2}-2)^{2}}{2}+2\nu_{12}^{2}-4\\ &=\frac{\nu_{12}(\nu_{12}-1)(6-\nu_{12}^{2})}{2}.\end{aligned}$ (5.51)
Here, we estimate $\nu_{12}$. First, we have
$0\leq\left|u_{1}+u_{2}+u_{3}\right|^{2}=3+2c_{12}+2c_{13}+2c_{23}.$ (5.52)
By (5.52) and $c_{12}=\max{(c_{12},c_{13},c_{23})}$, we have $\frac{1}{2}\leq c_{12}\leq 1$. Therefore, we have $1\leq\nu_{12}\leq 2$, and we obtain
$\begin{aligned} \frac{\nu_{12}(\nu_{12}-1)(6-\nu_{12}^{2})}{2}&\geq\nu_{12}-1\\ &=\frac{2c_{12}+1}{\sqrt{2(c_{12}+1)}+1}\\ &\geq\frac{2}{3}\left(c_{12}+\frac{1}{2}\right).\end{aligned}$ (5.53)
By (5.51) and (5.53), we obtain
@@
Hence (5.48) holds in the case $c_{23}\leq c_{13}\leq c_{12}$, and we complete the proof. ∎
Using this lemma, we estimate $c_{12},\ c_{13},\ c_{23}$. Before we estimate $c_{12},\ c_{13},\ c_{23}$, we define $\mathfrak{C}$ as
$\mathfrak{C}(t)=\left|c_{12}(t)+\frac{1}{2}\right|+\left|c_{23}(t)+\frac{1}{2}\right|+\left|c_{13}(t)+\frac{1}{2}\right|.$ (5.56)
@@
We have for $t\geq 0$,
$\frac{1}{4}\mathfrak{C}(t)\leq\max{\left(c_{12}(t)+\frac{1}{2},c_{13}(t)+\frac{1}{2},c_{23}(t)+\frac{1}{2}\right)}\leq\mathfrak{C}(t).$ (5.57)
@@
The right-hand inequality in (5.57) is obvious. It therefore suffices to prove the left-hand inequality. Since we have
$0\leq|u_{1}+u_{2}+u_{3}|^{2}=3+2c_{12}+2c_{13}+2c_{23},$
we have
$\left(c_{12}+\frac{1}{2}\right)+\left(c_{13}+\frac{1}{2}\right)+\left(c_{23}+\frac{1}{2}\right)\geq 0.$ (5.58)
By symmetry, we may assume that $c_{23}\leq c_{13}\leq c_{12}$. When $c_{23}\geq-\frac{1}{2}$, we have
$\mathfrak{C}=\left(c_{12}+\frac{1}{2}\right)+\left(c_{13}+\frac{1}{2}\right)+\left(c_{23}+\frac{1}{2}\right),$
and we obtain
$\frac{1}{3}\mathfrak{C}\leq\max{\left(c_{12}+\frac{1}{2},c_{13}+\frac{1}{2},c_{23}+\frac{1}{2}\right)}.$
When $c_{23}<-\frac{1}{2}\leq c_{13}$, by (5.58), we have
$2\left(c_{12}+\frac{1}{2}\right)$ $\geq\left(c_{12}+\frac{1}{2}\right)+\left(c_{13}+\frac{1}{2}\right)$
$\geq-\left(c_{23}+\frac{1}{2}\right)=\left|c_{23}+\frac{1}{2}\right|.$
Therefore, we obtain
$4\left(c_{12}+\frac{1}{2}\right)$ $\geq\left(c_{12}+\frac{1}{2}\right)+\left(c_{13}+\frac{1}{2}\right)+2\left(c_{12}+\frac{1}{2}\right)\geq\mathfrak{C}.$
When $c_{23}\leq c_{13}<-\frac{1}{2}\leq c_{12}$, by (5.58), we have
$\left(c_{12}+\frac{1}{2}\right)$ $\geq-\left(c_{13}+\frac{1}{2}\right)-\left(c_{23}+\frac{1}{2}\right)$
$=\left|c_{13}+\frac{1}{2}\right|+\left|c_{23}+\frac{1}{2}\right|.$
Thus, we have
$2\left(c_{12}+\frac{1}{2}\right)\geq\left(c_{12}+\frac{1}{2}\right)+\left|c_{13}+\frac{1}{2}\right|+\left|c_{23}+\frac{1}{2}\right|=\mathfrak{C}.$
Gathering these arguments, we obtain (5.57). ∎
Lemma 5.9.
Let $\vec{u}$ be a (1,3)-sign solitary waves. Then, we have for $t\gg 1$
$\left|c_{12}(t)+\frac{1}{2}\right|+\left|c_{23}(t)+\frac{1}{2}\right|+\left|c_{13}(t)+\frac{1}{2}\right|\lesssim\frac{1}{(\log{t})^{\frac{1}{2}}}.$ (5.59)
Proof.
First, by Lemma 5.6, there exists $C>0$ such that we have for $t\gg 1$,
@@
Here, we assume that there exists $t^{\prime}\gg 1$ such that
$\max{\left(c_{12}(t^{\prime})+\frac{1}{2},c_{13}(t^{\prime})+\frac{1}{2},c_{23}(t^{\prime})+\frac{1}{2}\right)}>\frac{45C}{L^{\frac{1}{2}}(t^{\prime})}.$ (5.63)
@@
$\max{\left(c_{12}(t)+\frac{1}{2},c_{13}(t)+\frac{1}{2},c_{23}(t)+\frac{1}{2}\right)}\geq\frac{45C}{L^{\frac{1}{2}}(t)}.$ (5.64)
@@
Since $L^{-\frac{1}{2}}$ is decreasing, (5.64) continues to hold for $0<t-T_{1}\ll 1$, contradicting the definition of $T_{1}$. Therefore $T_{1}=\infty$. In particular, the above computation shows that $\max(c_{12},c_{13},c_{23})$ is monotone increasing, and since it is bounded above by (5.62), it converges as $t\to\infty$ to some $c_{\max}\in\left(-\frac{1}{2},\frac{1}{3}\right]$. For $t$ sufficiently large, if $c_{12}$ is the maximum, then
@@
From this, taking $T_{2}\gg 1$ sufficiently large, we obtain
$1$ $\gtrsim\max{(c_{12}(e^{T_{2}}),c_{13}(e^{T_{2}}),c_{23}(e^{T_{2}}))}-\max{(c_{12}(T_{2}),c_{13}(T_{2}),c_{23}(T_{2}))}$
$\gtrsim\int_{T_{2}}^{e^{T_{2}}}\frac{ds}{s\log{s}},$
which is a contradiction. Hence we have for $t\gg 1$
$\max{\left(c_{12}(t)+\frac{1}{2},c_{13}(t)+\frac{1}{2},c_{23}(t)+\frac{1}{2}\right)}\lesssim\frac{1}{L^{\frac{1}{2}}(t)}\sim\frac{1}{(\log{t})^{\frac{1}{2}}}.$ (5.66)
By Lemma 5.8 and (5.66), we obtain (5.59). ∎
@@ 6.1. Sharp estimates for lengths and angles
In the previous section, we obtained (5.59). This allows us to refine the estimates for $\rho_{1},\rho_{2},\rho_{3}$.
@@
Let $\vec{u}$ be a (1,3)-sign solitary waves. Then, we have for $t\gg 1$
@@
We first obtain a joint estimate for $a_{1},a_{2},a_{3}$ and $d_{12},d_{13},d_{23}$.
Lemma 6.2.
Let $\vec{u}$ be a (1,3)-sign solitary waves. Then, we have for $t\gg 1$
$|a_{1}|+|a_{2}|+|a_{3}|+|d_{12}|+|d_{13}|+|d_{23}|\lesssim\frac{1}{(\log{t})^{3}}.$ (6.5)
Before we start the proof, we define
@@
$\begin{aligned} \left|a_{1}-\frac{1}{5}(4d_{12}+4d_{13}+2d_{23})\right|&\lesssim\frac{|d|}{(\log{t})^{\frac{1}{4}}}+\frac{1}{(\log{t})^{3}},\\ \left|a_{2}-\frac{1}{5}(4d_{12}+2d_{13}+4d_{23})\right|&\lesssim\frac{|d|}{(\log{t})^{\frac{1}{4}}}+\frac{1}{(\log{t})^{3}},\\ \left|a_{3}-\frac{1}{5}(2d_{12}+4d_{13}+4d_{23})\right|&\lesssim\frac{|d|}{(\log{t})^{\frac{1}{4}}}+\frac{1}{(\log{t})^{3}}.\end{aligned}$ (6.6)
We first refine the estimates for the differential equations satisfied by $a_{1},a_{2},a_{3}$. By (2.12) and (6.1), we have
$\left|\mathcal{F}(\rho_{k})-(1-a_{k})\mathcal{F}(L)\right|\lesssim\frac{\mathcal{F}(L)|a|}{L^{\frac{1}{2}}}\sim\frac{|a|}{t(\log{t})^{\frac{1}{2}}}.$ (6.7)
@@
$\quad+O\left(\frac{|a|}{t(\log{t})^{\frac{1}{2}}}+\frac{1}{t(\log{t})^{6}}\right),$
@@
$\quad+O\left(\frac{|a|}{t(\log{t})^{\frac{1}{2}}}+\frac{1}{t(\log{t})^{6}}\right).$
By the same calculation, $a_{1},a_{2},a_{3}$ satisfy
$\begin{aligned} \dot{a}_{1}&=\left(-2a_{1}+\frac{1}{2}a_{2}+\frac{1}{2}a_{3}+d_{12}+d_{13}\right)\mathcal{F}(L)\\ &\quad+O\left(\frac{|a|}{t(\log{t})^{\frac{1}{2}}}+\frac{1}{t(\log{t})^{6}}\right),\\ \dot{a}_{2}&=\left(\frac{1}{2}a_{1}-2a_{2}+\frac{1}{2}a_{3}+d_{12}+d_{23}\right)\mathcal{F}(L)\\ &\quad+O\left(\frac{|a|}{t(\log{t})^{\frac{1}{2}}}+\frac{1}{t(\log{t})^{6}}\right),\\ \dot{a}_{3}&=\left(\frac{1}{2}a_{1}+\frac{1}{2}a_{2}-2a_{3}+d_{13}+d_{23}\right)\mathcal{F}(L)\\ &\quad+O\left(\frac{|a|}{t(\log{t})^{\frac{1}{2}}}+\frac{1}{t(\log{t})^{6}}\right).\end{aligned}$ (6.8)
@@
$\quad+O\left(\frac{|a|}{t(\log{t})^{\frac{3}{2}}}+\frac{1}{t(\log{t})^{7}}\right).$
@@
$\begin{aligned} \dot{d}_{12}&=\left(3d_{12}+\frac{3}{2}d_{13}+\frac{3}{2}d_{23}-\frac{3}{4}a_{1}-\frac{3}{4}a_{2}+\frac{3}{2}a_{3}\right)\frac{\mathcal{F}(L)}{L}\\ &\quad+O\left(\frac{|a|+|d|}{t(\log{t})^{\frac{3}{2}}}+\frac{1}{t(\log{t})^{7}}\right),\\ \dot{d}_{13}&=\left(\frac{3}{2}d_{12}+3d_{13}+\frac{3}{2}d_{23}-\frac{3}{4}a_{1}+\frac{3}{2}a_{2}-\frac{3}{4}a_{3}\right)\frac{\mathcal{F}(L)}{L}\\ &\quad+O\left(\frac{|a|+|d|}{t(\log{t})^{\frac{3}{2}}}+\frac{1}{t(\log{t})^{7}}\right),\\ \dot{d}_{23}&=\left(\frac{3}{2}d_{12}+\frac{3}{2}d_{13}+3d_{23}+\frac{3}{2}a_{1}-\frac{3}{4}a_{2}-\frac{3}{4}a_{3}\right)\frac{\mathcal{F}(L)}{L}\\ &\quad+O\left(\frac{|a|+|d|}{t(\log{t})^{\frac{3}{2}}}+\frac{1}{t(\log{t})^{7}}\right).\end{aligned}$ (6.9)
@@
$|\dot{d}_{12}|+|\dot{d}_{13}|+|\dot{d}_{23}|\lesssim\frac{|a|+|d|}{t\log{t}}+\frac{1}{t(\log{t})^{7}}.$ (6.10)
@@
Then, by direct computation, we have
@@
$\begin{aligned} \dot{\xi}_{1}&=\left(-2\xi_{1}+\frac{1}{2}\xi_{2}+\frac{1}{2}\xi_{3}\right)\mathcal{F}(L)+O\left(\frac{|\xi|+|d|}{t(\log{t})^{\frac{1}{2}}}+\frac{1}{t(\log{t})^{6}}\right),\\ \dot{\xi}_{2}&=\left(\frac{1}{2}\xi_{1}-2\xi_{2}+\frac{1}{2}\xi_{3}\right)\mathcal{F}(L)+O\left(\frac{|\xi|+|d|}{t(\log{t})^{\frac{1}{2}}}+\frac{1}{t(\log{t})^{6}}\right),\\ \dot{\xi}_{3}&=\left(\frac{1}{2}\xi_{1}+\frac{1}{2}\xi_{2}-2\xi_{3}\right)\mathcal{F}(L)+O\left(\frac{|\xi|+|d|}{t(\log{t})^{\frac{1}{2}}}+\frac{1}{t(\log{t})^{6}}\right).\end{aligned}$ (6.13)
@@
$\quad+O\left(\frac{|\xi|^{2}+|d|^{2}}{t(\log{t})^{\frac{1}{2}}}+\frac{1}{t(\log{t})^{6}}\right).$
Since we have
$\xi_{1}^{2}+\xi_{2}^{2}+\xi_{3}^{2}\geq\xi_{1}\xi_{2}+\xi_{2}\xi_{3}+\xi_{3}\xi_{1},$
there exists $c_{1}>0$ such that we have for $t\gg 1$
$\frac{d}{dt}|\xi|^{2}\leq-\frac{c_{1}}{t}|\xi|^{2}+\frac{1}{c_{1}}\left(\frac{|d|^{2}}{t(\log{t})^{\frac{1}{2}}}+\frac{1}{t(\log{t})^{6}}\right).$ (6.14)
By Gronwall’s inequality and (6.10), we obtain
$|\xi|^{2}\lesssim\frac{|d|^{2}}{(\log{t})^{\frac{1}{2}}}+\frac{1}{(\log{t})^{6}},$
@@
By (6.6) and (6.9), there exists $c_{2}>0$ such that
$\begin{aligned} \left|\dot{d}_{12}-\left(\frac{12}{5}d_{12}+\frac{9}{5}d_{13}+\frac{9}{5}d_{23}\right)\frac{\mathcal{F}(L)}{L}\right|&\leq c_{2}\left(\frac{\mathcal{F}(L)}{L^{\frac{5}{4}}}|d|+\frac{\mathcal{F}(L)}{L^{4}}\right),\\ \left|\dot{d}_{13}-\left(\frac{9}{5}d_{12}+\frac{12}{5}d_{13}+\frac{9}{5}d_{23}\right)\frac{\mathcal{F}(L)}{L}\right|&\leq c_{2}\left(\frac{\mathcal{F}(L)}{L^{\frac{5}{4}}}|d|+\frac{\mathcal{F}(L)}{L^{4}}\right),\\ \left|\dot{d}_{23}-\left(\frac{9}{5}d_{12}+\frac{9}{5}d_{13}+\frac{12}{5}d_{23}\right)\frac{\mathcal{F}(L)}{L}\right|&\leq c_{2}\left(\frac{\mathcal{F}(L)}{L^{\frac{5}{4}}}|d|+\frac{\mathcal{F}(L)}{L^{4}}\right).\end{aligned}$ (6.15)
@@
Furthermore, by Lemma 5.8, we have
$\max{(d_{12},d_{13},d_{23})}\sim|d|.$ (6.17)
@@
$\max{(d_{12}(t^{\prime}),d_{13}(t^{\prime}),d_{23}(t^{\prime}))}>\frac{100c_{2}}{L^{3}(t^{\prime})}.$ (6.18)
@@
Furthermore, we assume that $T_{1}<\infty$. Then, we assume
@@
Then, by (6.15) and (6.17), we have at $t=T_{1}$
$\dot{d}_{12}$ $\geq\left(\frac{12}{5}d_{12}+\frac{9}{5}d_{13}+\frac{9}{5}d_{23}\right)\frac{\mathcal{F}(L)}{L}-c_{2}\left(\frac{\mathcal{F}(L)}{L^{\frac{5}{4}}}|d|+\frac{\mathcal{F}(L)}{L^{4}}\right)$
$\geq\frac{3}{5}\frac{\mathcal{F}(L)}{L}d_{12}-c_{2}\left(\frac{\mathcal{F}(L)}{L^{\frac{5}{4}}}|d|+\frac{\mathcal{F}(L)}{L^{4}}\right)$
$\geq\frac{1}{2}\frac{\mathcal{F}(L)}{L}d_{12}-c_{2}\frac{\mathcal{F}(L)}{L^{4}}$